OpenStudy (anonymous):

which ordered pair is a solution of this equation? -7x-5y=42. explain.

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

What are the ordered pairs? Do you have a list of them?

OpenStudy (anonymous):

you mean like choices?

Directrix (directrix):

I suppose so. All I know is that your question says "which ordered pair" and I don't see any ordered pairs.

OpenStudy (anonymous):

ohh yeah hold on

OpenStudy (anonymous):

idk theres - (-6,0 ) - (0,-5 ) - (-5,0) -(0,-6)

OpenStudy (anonymous):

there ^

OpenStudy (anonymous):

how do i figure that out though

OpenStudy (triciaal):

like the one earlier you have (x, y) substitute the values for x an y on the left hand side of the equation and simplify. if your answer is the same as on the right hand side then that pair (x, y) is a solution. do this for each group.

OpenStudy (triciaal):

i'll do the first one.

OpenStudy (anonymous):

ohh so i can solve the choices and see if it matches the graph or equation?

OpenStudy (triciaal):

-7x-5y=42 (-6, 0) x = -6 and y = 0 left hand side -7(-6)-5(0) = 42 -0 = 42 right hand side = 42 42 = 42 therefore (-6, 0) is a solution

OpenStudy (anonymous):

okay let me write these steps down so i wont forget and i can try some on my own

OpenStudy (anonymous):

i still dont understand the steps, like i need more explanation with steps

OpenStudy (mathmath333):

put the options in \(\large\tt \begin{align} \color{black}{ -7x-5y }\end{align}\) and see by which option u get \(42\)

OpenStudy (mathmath333):

example 1st option (-6,0) \(\large\tt \begin{align} \color{black}{ -7x-5y\\ =-7\times -6-5\times 0\\ =+42-0\\ =42 }\end{align}\) so first option is the ordered pair

OpenStudy (anonymous):

i need the steps first

OpenStudy (mathmath333):

^ steps

OpenStudy (anonymous):

oh i see

OpenStudy (mathmath333):



OpenStudy (mathmath333):

what is \(\Large -6\times -7\)

OpenStudy (anonymous):

idk :(

OpenStudy (mathmath333):

its \(\Large +42\)